# Understanding Language Prior of LVLMs by Contrasting Chain-of-Embedding

### v2

@@ Abstract
Large vision-language models (LVLMs) achieve strong performance on multimodal tasks, yet they often default to their language prior (LP)—memorized textual patterns from pre-training while under-utilizing visual evidence. Prior analyses of LP mostly rely on input–output probing, which fails to reveal the internal mechanisms governing when and how vision influences model behavior. To address this gap, we present the first systematic analysis of language prior through the lens of chain-of-embedding, which examines the layer-wise representation dynamics within LVLMs. Our analysis reveals a universal phenomenon: each model exhibits a Visual Integration Point (VIP), a critical layer at which visual information begins to meaningfully reshape hidden representations and influence decoding. Building on this observation, we introduce the Total Visual Integration (TVI) estimator, which aggregates representation distance beyond the VIP to quantify how strongly visual query influences response generation. Across 54 model–dataset combinations spanning 9 contemporary LVLMs and 6 benchmarks, we demonstrate that VIP consistently emerges, and that TVI reliably predicts the strength of language prior. This offers a principled toolkit for diagnosing and understanding language prior in LVLMs.               Code:
@@ 1 Introduction
Understanding and quantifying LP in LVLMs is thus critical, both for diagnosing their limitations and for guiding the design of more reliable multimodal systems. However, current approaches to analyzing LP primarily rely on input–output probing. For instance, Lee et al. (2025) and Luo et al. (2025) constructed datasets with counterfactual visual input to measure models’ performance under challenging visual grounding scenarios, while Deng et al. (2025) evaluate models on modality-conflicting queries to assess modality preference. While useful, such coarse input-output analyses have fundamental limitation to investigate LP of LVLMs in-depth, because: (1) they ignore the rich latent representations inside the model, which may inform how textual and visual signals are integrated, and (2) they cannot disentangle where in the model the LP begins to interfere with effective visual integration, leaving per-sample mechanistic interpretation (Bereska & Gavves, 2024) elusive.
@@
Motivated by this, we propose a new framework for understanding and quantifying language prior, which leverages the chain-of-embedding—the sequence of hidden representations across LVLM layers. Making use of these latent representations is essential, because they provide direct insight into the inner mechanisms of LP, beyond surface-level outputs. Specifically, our framework contrasts embeddings from vision–text inputs ($Z_{\text{vis}}^{l}$) with those from vision-removed inputs ($Z_{\text{blind}}^{l}$), at each layer $l$. Based on the contrastive chain-of-embedding, we reveal a striking phenomenon: LVLMs exhibit a Visual Integration Point (VIP), a layer at which visual information begins to meaningfully influence the LVLM’s decoding process. At and beyond VIP, the distance between $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ increases substantially for vision-dependent tasks, signaling that the model has begun to actively integrate visual evidence to solve the task. In contrast, vision-independent tasks show a smaller such shift. Thus, VIP captures a critical point where visual input begins to exert meaningful influence on inference, revealing the extent to which the model relies on vision or falls back on language priors.
@@
Inspired by observations from VIP, we propose quantifying LP through Total Visual Integration (TVI), which measures the effective amount of visual integration that affects the answer decoding of LVLM. Specifically, TVI aggregates distance between contrastive embeddings $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ across all post-VIP layers to measure the cumulative strength of visual integration. Intuitively, TVI is inversely related to the magnitude of LP: models with strong reliance on language priors exhibit low TVI, while those that leverage vision more deeply exhibit high TVI. Through extensive experiments covering 9 contemporary LVLMs and 6 datasets (54 settings combined), we show the universality of the existence of VIP, and that TVI can be a reliable indicator of LP. Moreover, we demonstrate that TVI strongly correlates with performance on benchmarks requiring visual reasoning, outperforming other proxies such as visual attention weights or output divergence. Then, we provide a theoretical interpretation of our VIP measurement as well as analytic bounds of it for broader use in practice. We illustrate the overall framework in Figure 1, and summarize our contribution as follows:
@@
    Across 9 representative LVLMs and 6 datasets, we consistently demonstrate the existence of VIP, show how we can use it to predict the strength of language prior of an LVLM on a certain sample through TVI, and further present theoretical analyses on our framework.
@@ 2 Problem Statement
An LVLM has a vast amount of knowledge in its parameters obtained during unimodal pretraining and visual instruction tuning of entire model components. Since the pre-training of LLM backbone is far more extensive in quantity and diversity of data, and total computing budget, LVLMs are prone to over-reliance on memorized statistical textual patterns without integrating visual information during inference. Given an input $x$ and an LVLM $F_{\theta}$, we define the model’s reliance on statistical textual patterns as the language prior, $\text{LP}(x,F_{\theta})$. Note that LP is more like a latent property that lacks a gold-standard measurement. Therefore, previous work typically approximates how robust an LVLM is against LP through its performance on carefully constructed datasets. In contrast, we propose a novel approach that (1) does not require any annotations or careful data curation, (2) tries to quantify LP in a more direct manner, which enables flexible and fine-grained, sample-wise diagnosis for LP of LVLMs. Refer to Appendix A for additional context.
@@
Although there have been recent attempts in analyzing LP in LVLMs, they primarily focus on evaluating model predictions over curated datasets (Lee et al., 2025; Luo et al., 2025; Vo et al., 2025), without offering a well-defined or generalizable formulation. We argue that such coarse input–output analysis is insufficient: it cannot reveal how LP manifests within the model nor how it can be rigorously quantified. In particular, prior approaches overlook the rich latent information encoded inside the LLM component of an LVLM—intermediate representations that may inform how visual and textual signals are integrated and how biases emerge. Making use of these latent representations is essential because they provide direct insight into the inner mechanisms of LP, beyond surface-level outputs. With this motivation, we pose the following research question:
Can we derive a formal framework to understand and quantify the language prior of LVLMs
through the lens of their internal states?
@@ 3.1 Chain-of-Embedding and Representation Distance
In contrast to previous approaches that focus on LVLM output (Rahmanzadehgervi et al., 2024; Vo et al., 2025; Lee et al., 2025; Luo et al., 2025), we leverage the chain-of-embedding for fine-grained analysis of LVLM, which is defined as a sequence of hidden states across layers, i.e., $(Z^{1},\cdots,Z^{L})$, where $Z^{l}=f_{l}(X_{v},X_{t})\in\mathbb{R}^{d_{z}}$ [^1] Although $Z^{l}=f_{l}(...f_{2}(f_{1}(X_{v},X_{t})))$ is more precise, we slightly abuse the notation for clarity. denotes a $d_{z}$-dimensional output token embedding at layer $l\in\{1,...,L\}$. Importantly, we contrast embeddings from two different input constructions as below.
@@
Now, given a distance metric $d$, and $\mathcal{D}_{\star}\sim\mathcal{P}_{\star}$, we analyze the difference between these two embeddings per layer by defining an expected representation distance and its finite-sample estimator,
$\mathbf{D}_{l}(\mathcal{P}_{\star},F_{\theta}):=\mathbb{E}_{(X_{v},X_{t})\sim\mathcal{P}_{\star}}[d(Z_{\text{vis}}^{l},Z_{\text{blind}}^{l})]\text{,}\quad\mathbf{D}_{l}(\mathcal{D}_{\star},F_{\theta}):=\frac{1}{|\mathcal{D}_{\star}|}\sum_{(x_{v},x_{t})_{i}\in\mathcal{D}_{\star}}d(z_{\text{vis}}^{l,i},z_{\text{blind}}^{l,i})\text{.}$ (2)
We adopt the cosine distance by default, though other distance functions, including non-metric distances (Deza & Deza, 2009), can also be valid. An ablation study with alternative metrics is provided in Section 4. Intuitively, $Z_{\text{vis}}^{l}$ should encode distinctive visual semantics that cannot be inferred from text alone, whereas $Z_{\text{blind}}^{l}$ primarily reflects the model’s default linguistic expectations. However, the degree of this discrimination can depend on how visual information contributes differently to different data, and across different layers $l$ of the model. We elaborate on this in the next section.
@@ 3.2 Visual Integration Point Hypothesis
Deep neural networks are known to develop hierarchical representations across layers (Chen et al., 2023; Fan et al., 2024; Jin et al., 2025), where each layer has different types and resolutions of information (Joseph & Nanda, 2024; Skean et al., 2025; Artzy & Schwartz, 2024; Jiang et al., 2025). In this paper, we hypothesize that an LVLM has a Visual Integration Point (VIP) $l^{*}$, a critical layer where the model begins to actively leverage visual information to perform task-specific reasoning. Prior to this point, the model primarily engages in general-purpose processing of visual and textual inputs—visual features may be “seen,” but not yet “used” to guide inference, and the interactions between modalities remain shallow. This behavioral shift can be reflected in the representation distances: at and beyond VIP, the distance between $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ increases markedly for vision-dependent tasks ($\mathcal{P}_{\text{TV}}$), signaling that the model has started to utilize visual information to solve the task, while vision-independent tasks ($\mathcal{P}_{\text{T}}$) show smaller such shift. Thus, the notion of VIP captures a key behavior transition inside LVLMs. If such a specific point $l^{*}$ exists, identifying it allows us to localize where the differences between language-prior-dominated and visually grounded inference start to manifest within the model’s internal processing. We formalize this hypothesis below.
@@
where $\tau\in\mathbb{R}^{+}$ denote a model-dependent constant threshold for each data distribution.
@@ 3.3 Quantifying Language Prior of LVLMs through Total Visual Integration
$\text{TVI}(l^{*};x,F_{\theta})=\frac{1}{L-l^{*}+1}\sum_{l=l^{*}}^{L}\big[d(z^{l}_{\text{vis}},z^{l}_{\text{blind}}))\big]\text{,}$ (4)
@@
Here $l^{*}$ marks the VIP layer, where visual information begins to meaningfully influence the model’s internal states. The TVI score then measures the cumulative contribution of visual information by averaging representation distances across all subsequent layers ($l\geq l^{*}$). The idea behind TVI is that once the model passes the visual integration point, its internal representations increasingly reflect meaningful visual grounding, rather than shallow alignment or language-driven statistical patterns. A higher TVI indicates that visual information exerts a stronger and more persistent influence throughout the later layers, while a lower TVI suggests that the model is more likely to remain text-dominated even after $l^{*}$. In this sense, TVI provides a holistic measure of how much the model truly uses vision for actual problem solving: a strong LP corresponds to weak or shallow visual integration (low TVI), while effective multimodal reasoning corresponds to high TVI.
@@
0.7174
($p<0.001$)
Taken together, results from Figure 2 and Table 1 highlight two key insights: (1) the existence of the visual integration point $l^{*}$, where effective problem-solving starts to happen by integrating visual information, and (2) the strong relationship between post-$l^{*}$ TVI and downstream performance on vision-dependent tasks. These findings demonstrate that VIP and TVI provide a principled toolkit for analyzing visual integration and language prior in LVLMs. We summarize our findings below.
@@ 4 Extended Experiments
We extend the experimental setups described in Section 3 to a broader range of 6 datasets and 9 LVLMs, including Qwen2.5-VL-7B (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), Gemma-3-4B (Team et al., 2025), LLaVA-v1.5-7B (Liu et al., 2024a), Eagle2.5-8B (Chen et al., 2025a), Llama-3.2-11B-Vision [^2] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision-Instruct, LLaVA-NeXT-Vicuna-7B (Liu et al., 2024b), LLaVA-OV-Qwen2-7B (Li et al., 2025) and SmolVLM (Marafioti et al., 2025). We use the instruction-tuned version for all models. For the dataset, we consider general VQA benchmarks including MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), MMStar (Chen et al., 2024), and MMMU (Yue et al., 2024). We also incorporate two benchmarks specifically designed for language prior evaluation, which are VLind-Bench (Lee et al., 2025) and ViLP (Luo et al., 2024). This results in a combination of 54 experimental settings. Implementation details, including data statistics, generation configuration, strategy for VIP selection, etc., are provided in Appendix C. As illustrated in Figure 3, the emergence of VIP is remarkably consistent across all settings: for each model, there exits a clear transition layer $l^{*}$ where the distance between embeddings $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ increases more significantly for vision-dependent group ($\mathcal{D}_{\text{TV}}$), compared to the vision-independent group ($\mathcal{D}_{\text{T}}$). These results highlight the universality of the VIP existence. The complete experimental results are provided in Appendix D.
@@
To examine whether TVI (c.f. Definition 3.2) reliably reflects the strength of the language prior, we contrast results on two complementary datasets: ViLP and MMBench. ViLP is intentionally constructed to induce a strong LP on the data side by designing queries where plausible answers can often be inferred from textual patterns or statistical correlations without the need for visual grounding. In contrast, MMBench represents a weak LP setting, with less misleading questions that elicit more visual grounding. As shown in Figure 4, our analysis reveals that datasets with stronger language priors (e.g., ViLP) yield lower TVI values, indicating weaker visual integration in the model, whereas less biased datasets (e.g., MMBench) produce higher TVI values, reflecting stronger use of visual information. This confirms that TVI serves as a reliable quantitative indicator of LP.
@@
0.7155
$(p<0.001)$
0.6335
$(p<0.001)$
0.6727
$(p<0.001)$
0.5709
$(p<0.001)$
Visual
Attention
0.0871
$(p=0.075)$
-0.0364
$(p=0.530)$
0.4967
$(p<0.001)$
0.0746
$(p=0.197)$
Output
Divergence
0.2978
$(p<0.001)$
0.5084
$(p<0.001)$
0.1627
$(p<0.001)$
0.5615
$(p<0.001)$
@@
Ablations on model scales.
Figure 5: Ablations on model scales. VIP and the dimension-normalized TVI analysis results for three variants of Gemma-3 model family.
We further examine whether our findings generalize across models of different scales. As shown in Figure 5, a clearly defined VIP consistently emerges across models of varying sizes (4B, 12B, and 27B), underscoring the robustness and generality of our proposed behavioral analysis framework. Interestingly, we also find that the VIP tends to appear at a similar relative depth, which is approximately 60% of the total number of layers, regardless of model size. In addition, after normalizing by the dimensionality of hidden states, we observe that the average normalized TVI is consistently higher in larger models on both $\mathcal{D}_{\text{VT}}$ and $\mathcal{D}_{\text{T}}$. This suggests that larger models are more effective at leveraging visual information in a uniform manner across diverse input types, thereby exhibiting greater robustness to misleading language priors. These observations collectively reinforce the broader applicability of our framework in analyzing visual integration behavior across model scales.

### v3

@@ Abstract
Large vision-language models (LVLMs) achieve strong performance on multimodal tasks, yet they often default to their language prior (LP)—memorized textual patterns from pre-training while under-utilizing visual evidence. Prior analyses of LP mostly rely on input–output probing, which fails to reveal the internal mechanisms governing when and how vision influences model behavior. To address this gap, we present the first systematic analysis of language prior through the lens of chain-of-embedding, which examines the layer-wise representation dynamics within LVLMs. Our analysis reveals a universal phenomenon: each model exhibits a Visual Integration Point (VIP), a critical layer at which visual information begins to meaningfully reshape hidden representations and influence decoding for multimodal reasoning. Building on this observation, we introduce the Total Visual Integration (TVI) estimator, which aggregates representational discrepancy beyond the VIP to quantify how strongly visual query influences response generation. Across 60 model–dataset combinations spanning 10 contemporary LVLMs and 6 benchmarks, we demonstrate that VIP consistently emerges, and that TVI reliably predicts the strength of language prior. This offers a principled toolkit for diagnosing and understanding language prior in LVLMs.       Code:
@@ 1 Introduction
Understanding and quantifying LP in LVLMs is thus critical, both for diagnosing their limitations and for guiding the design of more reliable multimodal systems. However, current approaches to analyzing LP primarily rely on input–output probing. For instance, Lee et al. (2025) and Luo et al. (2025) constructed datasets with counterfactual visual input to measure models’ performance under challenging visual grounding scenarios, while Deng et al. (2025) evaluate models on modality-conflicting queries to assess modality preference. While useful, such coarse input-output analyses have fundamental limitation to investigate LP of LVLMs in-depth, because: (1) they ignore the rich latent representations inside the model, which may inform how textual and visual signals are integrated, and (2) they cannot disentangle where in the model the LP begins to interfere with effective visual integration, leaving per-sample mechanistic interpretation (Bereska and Gavves, 2024) elusive.
@@
Motivated by this, we propose a new framework for understanding and quantifying language prior, which leverages the chain-of-embedding—the sequence of hidden representations across LVLM layers. Making use of these latent representations is essential, because they provide direct insight into the inner mechanisms of LP, beyond surface-level outputs. Specifically, our framework contrasts embeddings from vision-text inputs ($Z_{\text{vis}}^{l}$) with those from vision-removed inputs ($Z_{\text{blind}}^{l}$), at each layer $l$. Based on the contrastive chain-of-embedding, we reveal a striking phenomenon: LVLMs exhibit a Visual Integration Point (VIP), a layer at which visual information begins to meaningfully influence the LVLM’s decoding process. At and beyond VIP, the distance between $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ increases substantially for vision-dependent tasks, signaling that the model has begun to actively integrate visual evidence to solve the task. In contrast, vision-independent tasks show a smaller such shift. Thus, VIP captures a critical point where visual input begins to exert meaningful influence on inference, revealing the extent to which the model relies on vision or falls back on language priors.
@@
Inspired by observations from VIP, we propose quantifying LP through Total Visual Integration (TVI), which measures the effective amount of visual integration that affects the answer decoding of LVLM. Specifically, TVI aggregates distance between contrastive embeddings $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ across all post-VIP layers to measure the cumulative strength of visual integration. Intuitively, TVI is inversely related to the magnitude of LP: models with strong reliance on language priors exhibit low TVI, while those that leverage vision more deeply exhibit high TVI. Through extensive experiments covering 10 contemporary LVLMs and 6 datasets (60 settings combined), we show the universality of the existence of VIP, and that TVI can be a reliable indicator of LP. Moreover, we demonstrate that TVI strongly correlates with performance on benchmarks requiring visual reasoning, outperforming other proxies such as visual attention weights or output divergence. Then, we provide a theoretical interpretation of our measure as well as analytic bounds of it for broader use in practice. We illustrate the overall framework in Figure 1, and summarize our contribution as follows:
@@
    Across 10 representative LVLMs and 6 datasets, we consistently demonstrate the existence of VIP, show how we can use it to predict the strength of language prior of an LVLM on a certain sample through TVI, and further present theoretical analyses on our framework.
@@ 2 Problem Statement
An LVLM has a vast amount of knowledge in its parameters obtained during unimodal pretraining and visual instruction tuning of entire model components. Since the pre-training of LLM backbone is far more extensive in quantity and diversity of data, and total computing budget, LVLMs are prone to over-reliance on memorized statistical textual patterns without integrating visual information during inference. Given an input $x$ and an LVLM $F_{\theta}$, we define the model’s reliance on statistical textual patterns as the language prior, $\text{LP}(x,F_{\theta})$. Note that LP is more like a latent property that lacks a gold-standard measurement. Therefore, previous work typically approximates how robust an LVLM is against LP through its performance on carefully curated datasets. In contrast, we propose a novel approach that (1) does not require any annotations or careful data curation, (2) tries to quantify LP in a more direct manner, which enables flexible and fine-grained, sample-wise diagnosis for LP of LVLMs. Refer to Appendix A for additional context.
@@
Although there have been recent attempts to analyze LP in LVLMs, they primarily focus on evaluating model predictions on curated datasets (Lee et al., 2025; Luo et al., 2025; Vo et al., 2025), without offering a well-defined or generalizable formulation. We argue that such coarse input-output analysis is insufficient: it cannot reveal how LP manifests within the model nor how it can be rigorously quantified. In particular, prior approaches overlook the rich latent information encoded inside LVLM—intermediate representations that inform how visual and textual signals are integrated and how biases emerge. Making use of these latent representations is essential because they provide direct insight into the inner mechanisms of LP, beyond surface-level outputs. With this motivation, we pose the following research question: Can we derive a formal framework to understand and quantify the language prior of LVLMs through the lens of their internal states?
@@ 3.1 Chain-of-Embedding and Representation Distance
In contrast to previous approaches that focus on LVLM output (Rahmanzadehgervi et al., 2024; Vo et al., 2025; Lee et al., 2025; Luo et al., 2025), we leverage the chain-of-embedding for fine-grained analysis of LVLM, which is defined as a sequence of hidden states across layers, i.e., $(Z^{1},\cdots,Z^{L})$, where $Z^{l}=f_{l}(X_{v},X_{t})\in\mathbb{R}^{d_{z}}$ [^1] Although $Z^{l}=f_{l}(...f_{2}(f_{1}(X_{v},X_{t})))$ is more precise, we slightly abuse the notation for clarity. denotes the last-token embedding at $l\in\{1,...,L\}$ as a contextual summary vector [^2] Such last-token embeddings integrate information from all preceding tokens and are widely used to investigate model’s behavior when generating the next token (Jiang et al., 2024; Tian et al., 2024; Li et al., 2025b).. Notably, we contrast embeddings from two different input constructions as below.
@@
Now, given a distance metric $d$, we analyze the difference between these two embeddings per layer by defining an expected representation distance and its finite-sample estimator,
$\mathbf{D}_{l}(\mathcal{P}_{\star},F_{\theta}):=\mathbb{E}_{(X_{v},X_{t})\sim\mathcal{P}_{\star}}[d(Z_{\text{vis}}^{l},Z_{\text{blind}}^{l})]\text{,}\quad\mathbf{D}_{l}(\mathcal{D}_{\star},F_{\theta}):=\frac{1}{|\mathcal{D}_{\star}|}\sum_{(x_{v},x_{t})_{i}\in\mathcal{D}_{\star}}d(z_{\text{vis}}^{l,i},z_{\text{blind}}^{l,i}),$ (2)
where $\mathcal{D}_{\star}$ is $\mathcal{D}_{VT}$ or $\mathcal{D}_{T}$, and $\mathcal{P}_{\star}$ is $\mathcal{P}_{VT}$ or $\mathcal{P}_{T}$.
We adopt the cosine distance by default, though other distance functions, including non-metric distances (Deza and Deza, 2009), can also be valid. An ablation study with alternative metrics is provided in Section 4. Intuitively, $Z_{\text{vis}}^{l}$ should encode distinctive visual semantics that cannot be inferred from text alone, whereas $Z_{\text{blind}}^{l}$ primarily reflects the model’s default linguistic expectations. However, the degree of this discrimination can depend on how visual information contributes differently to different data, and across different layers $l$ of the model. We elaborate on this in the next section.
@@ 3.2 Visual Integration Point Hypothesis
Deep neural networks are known to develop hierarchical representations across layers (Chen et al., 2023; Fan et al., 2024; Jin et al., 2025), where each layer has different types and resolutions of information (Joseph and Nanda, 2024; Skean et al., 2025; Artzy and Schwartz, 2024; Jiang et al., 2025). In this paper, we hypothesize that an LVLM has a Visual Integration Point (VIP) $l^{*}$, a critical layer where the model begins to actively leverage visual information to perform task-specific reasoning. Prior to this point, the model primarily engages in general-purpose processing of visual and textual inputs—visual features may be “seen,” but not yet “used” to guide inference, and the interactions between modalities remain shallow. This behavioral shift can be reflected in the representation distances: at and beyond VIP, the distance between $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ increases markedly for vision-dependent tasks ($\mathcal{P}_{\text{TV}}$), signaling that the model has started to utilize visual information to solve the task, while vision-independent tasks ($\mathcal{P}_{\text{T}}$) show smaller such shift. Thus, the notion of VIP captures a key behavior transition inside LVLMs. If such a specific point $l^{*}$ exists, identifying it allows us to localize where the differences between language-prior-dominated and visually grounded inference start to manifest within the model’s internal processing. We formalize this hypothesis below.
@@
where $\tau\in\mathbb{R}^{+}$ denote a model-dependent constant threshold for each data distribution [^3] We manually select the $\tau$ and VIP for each model over the observed distances $\mathbf{D}_{l}$ for analysis convenience (see Appendix B and C for details on this manual selection and an automatic selection method as well)..
@@ 3.3 Quantifying Language Prior of LVLMs through Total Visual Integration
$\text{TVI}(l^{*};x,F_{\theta})=\frac{1}{L-l^{*}+1}\sum_{l=l^{*}}^{L}\big[d(z^{l}_{\text{vis}},z^{l}_{\text{blind}})\big]\text{,}$ (4)
@@
Here $l^{*}$ marks the VIP layer, where visual information begins to meaningfully influence the model’s internal states for visually-grounded decoding. The TVI score then measures the cumulative contribution of visual information by averaging representation distances across all subsequent layers ($l\geq l^{*}$). The idea behind TVI is that once the model passes the VIP, its internal representations increasingly reflect effective visual grounding, rather than shallow alignment or language-driven statistical patterns. A higher TVI indicates that visual information is more effectively utilized during the response decoding phase, while a lower TVI suggests that the model is more likely to remain text-dominated even after $l^{*}$. In this sense, TVI provides a holistic measure of how much the model truly uses vision for actual problem solving: a strong LP corresponds to weak or shallow visual integration (low TVI), while effective multimodal reasoning corresponds to high TVI.
@@
Taken together, results from Figure 2 and Table 1 highlight two key insights: (1) the existence of the visual integration point $l^{*}$, where effective representational shifts starts to happen by integrating visual information, and (2) the strong relationship between post-$l^{*}$ TVI and downstream performance on vision-dependent tasks. These findings demonstrate that VIP and TVI provide a principled toolkit for analyzing visual integration and language prior in LVLMs. We summarize our findings below.
@@ 4 Extended Experiments
We extend the experimental setups described in Section 3 to a broader range of 6 datasets and 10 LVLMs, including Qwen2.5-VL-7B (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), Gemma-3-4B (Team et al., 2025), LLaVA-v1.5-7B (Liu et al., 2024a), Eagle2.5-8B (Chen et al., 2025a), Llama-3.2-11B-Vision [^4] https://huggingface.co/meta-llama/Llama-3.2-11B-Vision-Instruct, LLaVA-NeXT-Vicuna-7B (Liu et al., 2024b), LLaVA-OV-Qwen2-7B (Li et al., 2025a) SmolVLM (Marafioti et al., 2025), and InstructBLIP-Vicuna-7B (Dai et al., 2023). For the datasets, we consider general VQA benchmarks including MME (Chaoyou et al., 2023), MMBench (Liu et al., 2024d), MMStar (Chen et al., 2024), and MMMU (Yue et al., 2024). We also incorporate two benchmarks specifically designed for language prior evaluation, which are VLind-Bench (Lee et al., 2025) and ViLP (Luo et al., 2025). This results in a combination of 60 experimental settings. Implementation details, including data statistics, generation configuration, strategy for VIP selection, etc., are provided in Appendix C. As illustrated in Figure 3, the emergence of VIP is remarkably consistent across all settings: for each model, there exits a clear transition layer $l^{*}$ where the distance between embeddings $Z_{\text{vis}}^{l}$ and $Z_{\text{blind}}^{l}$ increases more significantly for vision-dependent group ($\mathcal{D}_{\text{TV}}$), compared to the vision-independent group ($\mathcal{D}_{\text{T}}$). These results highlight the universality of the VIP existence. Due to the space limit, we defer the complete experimental results to Appendix D.
@@
To examine whether TVI (c.f. Definition 3.2) reliably reflects the strength of the language prior, we contrast results on two complementary datasets: ViLP and MMBench. ViLP is intentionally constructed to induce a strong LP on the data side by designing queries where plausible answers can often be inferred from textual patterns or statistical correlations without the need for visual grounding. In contrast, MMBench represents a weak LP setting, with less misleading questions that encourage stronger visual grounding for task success. As shown in Figure 4, our analysis reveals that datasets with stronger language priors (e.g., ViLP) yield lower TVI values, indicating weaker visual integration in the model, whereas less biased datasets (e.g., MMBench) produce higher TVI values, reflecting stronger use of visual information. This confirms that TVI serves as a reliable quantitative indicator of LP.
Interventional validation for TVI.
Table 2: Downstream performance and TVI before and after intervention.
Accuracy (%) TVI
Before intervention 50.00 0.038
After intervention 52.33 0.144
To further verify whether TVI robustly quantifies LP under different inference setups, we conduct a small interventional study. Specifically, we applied an attention-correction-based hallucination mitigation method PAI (Liu et al., 2024c) to Qwen2.5-VL-7B as an inference-time intervention, which promotes the model to pay more attention to visual features, implicitly increasing visual integration. As shown from the results in Table 2, the intervention not only improves task performance but also yields a substantial increase in TVI. This observation demonstrates that TVI faithfully reflects changes in the model’s degree of visual integration, thereby providing robust evidence that it is a reliable metric for quantifying LP.
@@
Visual Attention 0.0871 $(p=0.075)$ -0.0364 $(p=0.530)$ 0.4967 $(p<0.001)$ 0.0746 $(p=0.197)$
Output Divergence 0.2978 $(p<0.001)$ 0.5084 $(p<0.001)$ 0.1627 $(p<0.001)$ 0.5615 $(p<0.001)$
@@
Varying model scales.
Figure 5: Varying model scales. VIP and the dimension-normalized TVI analysis results for three variants of Gemma-3 model family.
We further examine whether our findings generalize across models of different scales. As shown in Figure 5, the VIP consistently emerges across models of varying sizes (4B, 12B, and 27B), underscoring the robustness and generality of our proposed behavioral analysis framework. Interestingly, we also find that the VIP tends to appear at a similar relative depth, which is approximately 60% of the total number of layers, regardless of model size. In addition, after normalizing by the dimensionality of hidden states, we observe that the average normalized TVI is consistently higher in larger models on both $\mathcal{D}_{\text{VT}}$ and $\mathcal{D}_{\text{T}}$. This suggests that larger models are more effective at leveraging visual information in a uniform manner across diverse input types, thereby exhibiting greater robustness to misleading language priors. These observations collectively reinforce the broader applicability of our framework in analyzing visual integration behavior across model scales.
Practical utility.
We also investigate the practical applicability of TVI here by providing a concrete example to make use of our findings to actually improve LVLMs. Specifically, we leverage TVI as an additional regularization term along with the vanilla cross-entropy loss for next-token prediction during instruction tuning. That is, given the input $x=(x_{v},x_{t})$ and instruction $y$, we adjusted the original LLaVA training objective (Liu et al., 2023) to:
$\mathcal{L}(x,y;\theta)=-\log F_{\theta}(y|x)-\lambda\cdot\text{TVI}(l^{*};x,F_{\theta})\text{,}$ (5)
where the strength of the regularization is controlled by $\lambda$ (we set as 0.03). Due to resource limitations, we trained the model on a 60k randomly sampled subset of llava_v1_5_mix665k. All other hyperparameters used for visual instruction tuning remain the default of Liu et al. (2023).
Table 5: Effect of TVI regularization on downstream performance, MME dataset (Chaoyou et al., 2023) with LLaVA-v1.5-7B.
Perception Reasoning
LLaVA-v1.5 1369.75 298.21
LLaVA-v1.5 w/ TVI 1400.44 321.43
The result is shown in Table 5, where the performance improvement indicates that explicitly encouraging stronger visual integration (via TVI) leads to better downstream task performance. This highlights TVI’s potential as a helpful training regularizer for improving the visual perception and reasoning of LVLMs in practice.
Additional empirical analyses.
In addition, we deepen our understanding of the proposed framework by providing additional analysis in Appendix E, including different aggregation strategies for TVI calculation, instruction-level perturbation, image-text vs. image-only chain comparison, VIP and TVI evolution across training stages, and case studies on TVI failure cases.